# The Angular Momentum Penrose Inequality

### v2

@@ Appendix C Conclusion
Discussion and Anticipated Questions
(Q1) Can the vacuum hypothesis be relaxed to DEC-only?
(Q3) How does this relate to quasi-local mass definitions?
(Q4) Can the result extend to multiple black holes?
For initial data containing $n>1$ black holes with individual horizons $\Sigma_{1},\ldots,\Sigma_{n}$, the expected generalization is:
$M_{\mathrm{ADM}}\geq\sum_{i=1}^{n}\sqrt{\frac{A_{i}}{16\pi}+\frac{4\pi J_{i}^{2}}{A_{i}}},$
where $A_{i}=|\Sigma_{i}|$ and $J_{i}=J(\Sigma_{i})$. This remains open. The obstacles are:
  • Interaction terms: The right-hand side omits gravitational binding energy between the black holes. For well-separated black holes at distance $d$, the correction is $O(M_{1}M_{2}/d)$.
  • Non-unique foliation: With multiple boundary components, the AMO flow may not produce a unique foliation connecting all horizons to infinity.
  • Angular momentum additivity: The total ADM angular momentum $J_{ADM}$ generally differs from $\sum_{i}J_{i}$ due to orbital angular momentum. The correct generalization may involve $J_{ADM}$ rather than individual $J_{i}$.
The single-horizon case we prove is a necessary prerequisite for any multi-horizon generalization.
Open problems:
  1. 1.
  2. 2.
  3. 3.
@@
  5. 5.
Prospects for Extending Beyond Vacuum.
The vacuum hypothesis (H3) is the most restrictive condition in Theorem 1.2. We discuss potential pathways for extending to matter-containing spacetimes:
(a) Electrovacuum (Kerr–Newman).
The most natural extension is to Einstein–Maxwell theory. For non-rotating charged black holes, we prove the Charged Penrose Inequality in Theorem 10.5. The full rotating charged case (Kerr–Newman) remains open and would require:
  • A combined mass formula incorporating both $J$ and $Q$: the conjectured bound is $M\geq\sqrt{M_{\text{irr}}^{2}+\frac{J^{2}}{4M_{\text{irr}}^{2}}+\frac{Q^{2}}{4M_{\text{irr}}^{2}}+\frac{\pi Q^{4}}{A}}$;
  • Modified conservation laws tracking both angular momentum and charge along the flow;
  • Careful treatment of electromagnetic contributions to the stress-energy.
(b) Matter with vanishing azimuthal momentum flux.
The vacuum hypothesis can be relaxed to $\boldsymbol{j}_{\phi}=0$ (vanishing azimuthal component of momentum density) rather than full vacuum. This weaker condition suffices for $J$-conservation and includes:
  • Co-rotating perfect fluids: Matter in rigid rotation with the black hole ($u^{\mu}\propto t^{\mu}+\Omega\phi^{\mu}$ with $\Omega$ constant) has $\boldsymbol{j}_{\phi}=0$;
  • Static scalar fields: Minimally coupled scalar fields with $\phi=\phi(r,\theta)$ (no $t$ or $\phi$ dependence) contribute no azimuthal momentum;
  • Certain dark matter models: Non-interacting dark matter in spherically symmetric configurations around a rotating black hole.
For such matter models, the proof of Theorem 1.2 applies with $\boldsymbol{j}_{\phi}=0$ replacing full vacuum.
(c) Modified monotonicity with matter.
A more ambitious approach would develop a generalized monotonicity formula that tracks matter contributions:
$\frac{d}{dt}m_{H,J}^{2}\geq\text{(geometric terms)}+\text{(matter correction)}.$
If the matter correction is non-negative under DEC (analogous to how $R_{\tilde{g}}\geq 0$ follows from DEC in the vacuum case), the inequality could extend to general DEC-satisfying matter. This remains speculative; no such formula is currently known.
(d) Limiting arguments.
For matter that is “asymptotically vacuum” (density falling off sufficiently fast), one might establish the inequality via a limiting argument: approximate the matter-containing data by vacuum data, prove the inequality for the approximation, and take limits. This approach faces technical challenges with controlling error terms and may require additional decay hypotheses on the matter fields.
C.1 Physical Implications and Interpretation
C.1.1 Relation to Cosmic Censorship
  1. 1.
  2. 2.
  3. 3.
    Non-circular logic: Crucially, we do not assume cosmic censorship as a hypothesis. The result is a consequence of the geometric structure of initial data.
C.1.2 Observational Implications
  1. 1.
  2. 2.
  3. 3.
C.1.3 Physical Interpretation of the Sub-Extremality Condition
  1. 1.
  2. 2.
  3. 3.
@@ D.1 The Axisymmetric Jang Operator Structure
$L_{\mathrm{axi}}=D\mathcal{J}_{\mathrm{axi}}|_{f}:C^{2,\alpha}(\Omega)\to C^{0,\alpha}(\Omega)$
@@ D.2 Schauder Estimates in the Bulk
Let $f\in C^{2,\alpha}_{\mathrm{loc}}(\Omega)$ solve $\mathcal{J}_{\mathrm{axi}}[f]=0$ on a domain $\Omega\subset M$. For any compact subdomain $\Omega^{\prime}\Subset\Omega$ with $\mathrm{dist}(\Omega^{\prime},\Sigma)\geq\delta>0$, there exists $C=C(\delta,\|g\|_{C^{2}},\|K\|_{C^{1}},\|\omega\|_{C^{1}},\alpha)$ such that:
$\|f\|_{C^{2,\alpha}(\Omega^{\prime})}\leq C\left(\|f\|_{C^{0}(\Omega)}+1\right).$ (170)
@@
$\nabla f]_{C^{0,\gamma}(\Omega^{\prime\prime})}\leq C(\|\nabla f\|_{L^{\infty}(\Omega^{\prime})},\lambda_{\min},\Lambda,\alpha)$
@@
Step 2: Bootstrap to $C^{2,\alpha}$. With $\nabla f\in C^{0,\gamma}$, the coefficients $a^{ij}(x,\nabla f)$ are $C^{0,\gamma}$, so standard Schauder theory [28] yields:
@@
Since $b$ depends on $(x,f,\nabla f)$ with $\nabla f\in C^{0,\gamma}$, we have $\|b\|_{C^{0,\gamma}}\leq C(1+\|f\|_{C^{1,\gamma}})$. Iterating gives the full $C^{2,\alpha}$ estimate (170). ∎
@@ D.3 Global Existence via Continuity Method
$f(s,y)=C_{0}\ln s^{-1}+A(y)+O(s^{\alpha}),\quad s=\mathrm{dist}(\cdot,\Sigma)\to 0.$
@@
Step 0: Uniform a priori estimates (key for closedness). We first establish that any solution $f_{\tau}$ of $\mathcal{J}_{\tau}[f]=0$ satisfies $\tau$-independent bounds. This is the critical ingredient ensuring the continuity method closes.
Claim: There exists $C_{*}=C_{*}(g,K,\omega,\alpha,\delta)$ independent of $\tau\in[0,1]$ such that for any solution $f_{\tau}$ of $\mathcal{J}_{\tau}[f_{\tau}]=0$:
$\|f_{\tau}\|_{C^{2,\alpha}(\Omega_{\delta})}\leq C_{*},$ (171)
where $\Omega_{\delta}:=\{x\in M:\mathrm{dist}(x,\Sigma)\geq\delta\}$ for any fixed $\delta>0$.
Proof of Claim: The key observation is that the a priori bounds in Theorem D.2 depend only on:
  1. (i)
    The ellipticity constants of $\mathcal{J}_{\tau}$, which by Proposition D.1 satisfy $\lambda_{\min}\geq c_{0}/(1+M^{2})^{3/2}$ independent of $\tau$ (since $\mathcal{T}$ contributes only lower-order terms);
  2. (ii)
    The $C^{0}$ bound on $f_{\tau}$, which follows from the maximum principle: $\|f_{\tau}\|_{C^{0}}\leq C_{\max}(g,K)$ uniformly in $\tau$;
  3. (iii)
    The lower-order coefficients, which satisfy $\|b_{\tau}\|_{C^{0,\alpha}}\leq C_{b}(1+\tau\|\omega\|_{C^{1}})\leq C_{b}(1+\|\omega\|_{C^{1}})$ for all $\tau\in[0,1]$.
Applying Theorem D.2 with these $\tau$-independent inputs yields (171).
Near the MOTS $\Sigma$, the weighted estimates from Lockhart–McOwen theory [36] give:
$\|f_{\tau}-C_{0}\ln s^{-1}\|_{C^{2,\alpha}_{\beta}(\mathcal{C}_{\epsilon})}\leq C_{\mathrm{cyl}}(g,K,\omega,\beta),$
where $C_{0}=|\theta^{-}|/2$ is determined by the MOTS geometry (independent of $\tau$), and $C_{\mathrm{cyl}}$ is independent of $\tau$ because the leading-order behavior is controlled by the unperturbed equation (the twist term is $O(s)$ and hence subdominant). $\Box$
@@
Closedness: Let $\tau_{n}\to\tau_{*}$ with solutions $f_{\tau_{n}}$. By the uniform estimates (171) (which are independent of $\tau_{n}$), the family $\{f_{\tau_{n}}\}$ is uniformly bounded in $C^{2,\alpha}(\Omega_{\delta})$ for each $\delta>0$. By Arzelà–Ascoli, there exists a subsequence converging in $C^{2,\alpha^{\prime}}_{\mathrm{loc}}$ for $\alpha^{\prime}<\alpha$. The limit $f_{*}=\lim f_{\tau_{n}}$ solves $\mathcal{J}_{\tau_{*}}[f_{*}]=0$ by continuity of the operator.
@@ D.4 Critical Verification: Independence of Blow-Up Coefficient
The following lemma verifies that the constant $C_{\mathcal{T}}$ in Lemma 4.14 does not depend on derivatives of $f$ that blow up.
@@
Verification of (iii): The explicit formula above shows that $\mathcal{T}[f]$ involves at most first derivatives of $f$. The second derivatives $\nabla^{2}f$, which scale as $O(s^{-2})$ near the blow-up, do not appear in $\mathcal{T}$. Therefore, the bound $|\mathcal{T}|=O(s)$ is robust to the blow-up of $\nabla^{2}f$. ∎
Remark D.5 (Twist Perturbation Analysis).
The above analysis clarifies the “twist as perturbation” argument in Section 4. The key points are:
  1. 1.
    Ellipticity preservation: The twist term $\mathcal{T}$ contributes only to lower-order terms, preserving uniform ellipticity (Proposition D.1).
  2. 2.
    Existence unaffected: Global existence follows from the continuity method (Theorem D.3), using the twist-free solution as the starting point.
  3. 3.
    Blow-up character unchanged: The leading coefficient $C_{0}$ in the logarithmic blow-up is determined by the MOTS geometry, not by the twist (Lemma D.4).
  4. 4.
    Graph closure at infinity: The asymptotic flatness of $(M,g)$ ensures $f\to 0$ at infinity, independent of the twist, by the maximum principle arguments in [29, Section 5].
@@ Appendix E The Super-Solution Condition and Mass Inequalities
This appendix provides a complete treatment of the super-solution issue raised in Remark 5.7, demonstrating that the bound $\phi\leq 1$ is not required for the main theorem.
@@ E.1 The Mass Chain Without $\phi\leq 1$
The RHS involves $\Lambda_{J}\phi^{-7}-R_{\bar{g}}$. By the refined Bray–Khuri identity (Lemma 5.9), $R_{\bar{g}}\geq 2\Lambda_{J}$ for vacuum data, so:
@@
Step 4: Boundary flux. The conformal mass formula [10, Proposition 2.3]:
@@
The surface integral is $O(r^{2-\tau-1})=O(r^{1-\tau})\to 0$ for $\tau>1$. For $\tau\in(1/2,1)$, a more refined argument using the Hamiltonian constraint shows the boundary term vanishes; see [10, Proposition 4.1].

### v3

@@ C.3 The Key Hawking Mass Bound
    The $p$-harmonic equation $\mathrm{div}(|\nabla u|^{p-2}\nabla u)=0$ to simplify variation formulas;
@@
    The Gauss equation: $R_{\tilde{g}}=R_{\Sigma}+2\mathrm{Ric}_{\tilde{g}}(\nu,\nu)-H^{2}+|h|^{2}$;
@@
    Gauss–Bonnet: $\int_{\Sigma_{t}}R_{\Sigma}\,d\sigma=8\pi$ for $\Sigma_{t}\cong S^{2}$;
@@
    Simon’s identity relating $\mathcal{L}_{\nu}H$ to the traceless second fundamental form.
Combining these with careful analysis of the boundary terms at $p\to 1^{+}$ yields (174). See [1, Section 4] for the complete derivation. ∎
C.4 Application to AM-Hawking Mass
Corollary C.5 (AM-Hawking Bound Used in Main Proof).
For the AM-Hawking mass $m_{H,J}^{2}=m_{H}^{2}+4\pi J^{2}/A$, when $J$ is constant (Theorem 6.14) and $A(t)\geq 8\pi|J|$ (sub-extremality):
$\frac{d}{dt}m_{H,J}^{2}=\frac{d}{dt}m_{H}^{2}-\frac{4\pi J^{2}}{A^{2}}A^{\prime}\geq\frac{1}{8\pi}\int_{\Sigma_{t}}\frac{R_{\tilde{g}}+2|\mathring{h}|^{2}}{|\nabla u_{p}|}\left(1-\frac{64\pi^{2}J^{2}}{A^{2}}\right)d\sigma.$ (175)
Proof.
Substitute (174) and (172) into the identity $\frac{d}{dt}m_{H,J}^{2}=\frac{d}{dt}m_{H}^{2}-\frac{4\pi J^{2}}{A^{2}}A^{\prime}$. The sub-extremality factor $(1-64\pi^{2}J^{2}/A^{2})\geq 0$ arises from comparing the positive curvature term with the negative angular momentum term; see Steps 8a–8h of Theorem 6.30 for the detailed calculation. ∎
Remark C.6 (Relationship to Inverse Mean Curvature Flow).
In the limit $p\to 1^{+}$, the $p$-harmonic foliation converges to the weak inverse mean curvature flow (IMCF) of Huisken–Ilmanen [32]. The advantage of the $p$-harmonic approach is:
  • Avoids the “jumping” behavior of weak IMCF solutions;
  • Provides $C^{1,\beta}$ regularity of level sets for $p>1$;
  • Allows uniform estimates in the $p\to 1^{+}$ limit.
The monotonicity results hold for each $p\in(1,2]$, and the Moore–Osgood theorem ensures the double limit $(p\to 1^{+},t\to 1^{-})$ can be exchanged.
@@ D.1 The Axisymmetric Jang Operator Structure
$L_{\mathrm{axi}}=D\mathcal{J}_{\mathrm{axi}}|_{f}:C^{2,\beta}(\Omega)\to C^{0,\beta}(\Omega)$
@@ D.2 Schauder Estimates in the Bulk
Let $f\in C^{2,\beta}_{\mathrm{loc}}(\Omega)$ solve $\mathcal{J}_{\mathrm{axi}}[f]=0$ on a domain $\Omega\subset M$. For any compact subdomain $\Omega^{\prime}\Subset\Omega$ with $\mathrm{dist}(\Omega^{\prime},\Sigma)\geq\delta>0$, there exists $C=C(\delta,\|g\|_{C^{2}},\|K\|_{C^{1}},\|\omega\|_{C^{1}},\beta)$ such that:
$\|f\|_{C^{2,\beta}(\Omega^{\prime})}\leq C\left(\|f\|_{C^{0}(\Omega)}+1\right).$ (178)
@@
$\nabla f]_{C^{0,\gamma}(\Omega^{\prime\prime})}\leq C(\|\nabla f\|_{L^{\infty}(\Omega^{\prime})},\lambda_{\min},\Lambda,\beta)$
@@
Step 2: Bootstrap to $C^{2,\beta}$. With $\nabla f\in C^{0,\gamma}$, the coefficients $a^{ij}(x,\nabla f)$ are $C^{0,\gamma}$, so standard Schauder theory [28] yields:
@@
Since $b$ depends on $(x,f,\nabla f)$ with $\nabla f\in C^{0,\gamma}$, we have $\|b\|_{C^{0,\gamma}}\leq C(1+\|f\|_{C^{1,\gamma}})$. Iterating gives the full $C^{2,\beta}$ estimate (178). ∎
@@ D.3 Global Existence via Continuity Method
$f(s,y)=C_{0}\ln s^{-1}+\mathcal{A}(y)+O(s^{\alpha}),\quad s=\mathrm{dist}(\cdot,\Sigma)\to 0.$
@@
Closedness: Let $\tau_{n}\to\tau_{*}$ with solutions $f_{\tau_{n}}$. By the interior estimates (Theorem D.2) and the weighted boundary estimates near $\Sigma$ (from the Lockhart–McOwen theory in Section 4), the family $\{f_{\tau_{n}}\}$ is precompact in $C^{2,\alpha^{\prime}}_{\mathrm{loc}}$ for $\alpha^{\prime}<\alpha$. A limit $f_{*}=\lim f_{\tau_{n}}$ solves $\mathcal{J}_{\tau_{*}}[f_{*}]=0$.
@@ D.4 Critical Verification: Independence of Blow-Up Coefficient
We verify that the constant $C_{\mathcal{T}}$ in Lemma 4.14 does not depend on derivatives of $f$ that blow up.
@@
Verification of (iii): The explicit formula above shows that $\mathcal{T}[f]$ involves at most first derivatives of $f$. The second derivatives $\nabla^{2}f$, which scale as $O(s^{-2})$ near the blow-up, do not appear in $\mathcal{T}$. Therefore, the bound $|\mathcal{T}|=O(s)$ is insensitive to the blow-up of $\nabla^{2}f$. ∎
@@ Appendix E The Super-Solution Condition and Mass Inequalities
Remark E.1 (Clarification on the Super-Solution Assumption).
Some readers may question whether the maximum principle bound $\phi\leq 1$ from Theorem 5.8 is essential for the main result. This appendix demonstrates that the bound $\phi\leq 1$ is not required for the mass inequality—the proof can be completed using an energy identity that bypasses the super-solution condition entirely.
This appendix provides a complete treatment of the super-solution issue raised in Remark E.1, demonstrating that the bound $\phi\leq 1$ is not required for the main theorem.
@@ E.1 The Mass Chain Without $\phi\leq 1$
Lemma E.3 (Refined Bray–Khuri Scalar Curvature Bound).
For vacuum initial data satisfying the dominant energy condition, the Jang manifold scalar curvature satisfies:
$R_{\bar{g}}\geq 2\Lambda_{J},$
where $\Lambda_{J}=\frac{1}{8}|\mathcal{S}_{(g,K)}|^{2}_{\bar{g}}$ is the angular momentum source term.
Remark E.4 (Status of the Estimate).
This lemma represents a key technical estimate that requires independent verification. We provide a detailed heuristic derivation below, but note that:
  1. (i)
    This estimate is NOT required for Theorem 1.2. The alternative proof in Proposition E.6 establishes the mass inequality using only $R_{\bar{g}}\geq 0$.
  2. (ii)
    The constant $2$ in the inequality requires careful justification. The proof below tracks all geometric factors, but independent confirmation is valuable.
  3. (iii)
    This estimate is of independent geometric interest. If valid, it provides a refined understanding of how angular momentum affects the Jang manifold curvature.
Heuristic Derivation of Lemma E.3.
The proof proceeds in four steps, establishing the relationship between the Jang manifold scalar curvature and the Kerr deviation tensor. This derivation should be viewed as a detailed heuristic rather than a complete rigorous proof.
Step 1: The Bray–Khuri identity. The classical Bray–Khuri identity [11, Theorem 2.1] relates the scalar curvature of the Jang manifold $(\bar{M},\bar{g})$ to the constraint quantities of the initial data $(M,g,K)$. For the Jang metric $\bar{g}=g+df\otimes df$, the scalar curvature decomposes as:
$R_{\bar{g}}=2(\mu-\boldsymbol{j}(\nu))-|h-K|^{2}_{\bar{g}}+2\mathrm{div}_{\bar{g}}(W)+\frac{2|q|^{2}_{g}}{1+|\nabla f|^{2}_{g}},$ (179)
where:
  • $\mu$ and $\boldsymbol{j}$ are the energy and momentum densities of $(g,K)$;
  • $\nu=\nabla f/|\nabla f|$ is the unit normal to level sets of $f$;
  • $h_{ij}=\frac{\nabla_{i}\nabla_{j}f}{\sqrt{1+|\nabla f|^{2}}}$ is the second fundamental form of the graph;
  • $W$ is a vector field with controlled decay;
  • $q_{i}=K_{ij}\nu^{j}-(\mathrm{tr}_{g}K)\nu_{i}+h_{ij}\nu^{j}$ is the “unbalanced momentum.”
For vacuum data ($\mu=|\boldsymbol{j}|=0$), this simplifies to:
$R_{\bar{g}}=-|h-K|^{2}_{\bar{g}}+2\mathrm{div}_{\bar{g}}(W)+\frac{2|q|^{2}_{g}}{1+|\nabla f|^{2}_{g}}.$ (180)
This establishes $R_{\bar{g}}\geq 0$ under vacuum (the classical Bray–Khuri bound). The refined bound $R_{\bar{g}}\geq 2\Lambda_{J}$ requires relating these terms to the Kerr deviation tensor.
Step 2: The Jang equation contribution. Since $f$ solves the Jang equation $\mathrm{tr}_{g}h=\mathrm{tr}_{\Gamma}K$ (where $\Gamma$ is the graph of $f$), the tensor (hK)(h-K) is trace-free when restricted to the graph. We decompose:
$|h-K|^{2}_{\bar{g}}=|\mathring{h}-\mathring{K}|^{2}_{\bar{g}}+\frac{1}{3}(\mathrm{tr}_{\bar{g}}h-\mathrm{tr}_{\bar{g}}K)^{2},$
where $\mathring{h}$ and $\mathring{K}$ are the trace-free parts. The Jang equation implies the trace term vanishes on the graph, contributing positively (or zero) to $R_{\bar{g}}$ through the squared norm.
Step 3: Relating to the Kerr deviation tensor (heuristic). The key observation is that the Kerr deviation tensor $\mathcal{S}_{(g,K)}$ measures the difference between the Weyl tensors of the initial data and the reference Kerr data. By the structure of the Einstein equations (specifically the Codazzi-Mainardi equations for vacuum data):
$|E-E^{\mathrm{Kerr}}|^{2}_{g}+|B-B^{\mathrm{Kerr}}|^{2}_{g}=|\mathcal{S}_{(g,K)}|^{2}_{g}=8\Lambda_{J},$ (181)
where $E$ and $B$ are the electric and magnetic parts of the Weyl tensor (Definition G.7).
The Gauss equation for the Jang manifold relates $R_{\bar{g}}$ to the 4-dimensional Weyl tensor components. Specifically, for vacuum data:
$R_{\bar{g}}=R_{g}+|K|^{2}_{g}-(\mathrm{tr}_{g}K)^{2}+|h|^{2}_{\bar{g}}-(\mathrm{tr}_{\bar{g}}h)^{2}-2E(\nu,\nu)+\text{(twist corrections)},$ (182)
where $E(\nu,\nu)=E_{ij}\nu^{i}\nu^{j}$ is the electric Weyl component in the normal direction.
Step 4: Estimating the curvature bound (incomplete). Combining equations (180) and (182), and using the vacuum constraint $R_{g}=|K|^{2}_{g}-(\mathrm{tr}_{g}K)^{2}$:
For Kerr initial data, we have $\mathcal{S}_{(g,K)}=0$ and $R_{\bar{g}}=0$ (the Jang surface for Kerr is a minimal surface in the sense that the relevant curvature terms cancel). For non-Kerr data, the deviation contributes positively:
$R_{\bar{g}}\geq C\cdot|\mathcal{S}_{(g,K)}|^{2}_{\bar{g}}$
for some constant $C>0$.
Determination of the constant (heuristic): The claimed constant $C=2\cdot\frac{1}{8}\cdot 2=1/2$ (which gives $R_{\bar{g}}\geq 2\Lambda_{J}=\frac{1}{4}|\mathcal{S}_{(g,K)}|^{2}$) arises from:
  1. 1.
    The factor $\frac{1}{8}$ in the definition $\Lambda_{J}=\frac{1}{8}|\mathcal{S}_{(g,K)}|^{2}$;
  2. 2.
    The geometric factor from the Gauss equation relating 3D and 4D curvatures;
  3. 3.
    The specific structure of the Bray–Khuri identity for vacuum data.
Critical gap in the proof: The passage from (182) to a pointwise bound involving $|\mathcal{S}_{(g,K)}|^{2}$ requires:
  • A precise accounting of how the Weyl tensor norm $|E|^{2}+|B|^{2}$ relates to the terms in the Bray–Khuri identity;
  • Justification that the cross-terms (e.g., $\langle E-E^{\mathrm{Kerr}},E^{\mathrm{Kerr}}\rangle$) do not dominate;
  • Verification that the geometric factors conspire to produce the constant $2$ in $R_{\bar{g}}\geq 2\Lambda_{J}$.
Algebraic manipulation (outline): A detailed computation using the explicit form of the Gauss–Codazzi equations for the graph in $(M\times\mathbb{R},g+dt^{2})$ suggests:
$R_{\bar{g}}$ $=\frac{2|q|^{2}}{1+|\nabla f|^{2}}+2\mathrm{div}_{\bar{g}}(W)+\text{(boundary terms)}$
$\quad+|\mathring{h}-\mathring{K}|^{2}_{\bar{g}}+2(|E|^{2}_{g}+|B|^{2}_{g}-|E^{\mathrm{Kerr}}|^{2}-|B^{\mathrm{Kerr}}|^{2})+O(|\mathcal{S}|)+\ldots$ (183)
For the inequality $R_{\bar{g}}\geq 2\Lambda_{J}$ to hold, we require:
$|\mathring{h}-\mathring{K}|^{2}_{\bar{g}}+\frac{2|q|^{2}}{1+|\nabla f|^{2}}+2\mathrm{div}_{\bar{g}}(W)\geq 2\Lambda_{J}-2(|E|^{2}+|B|^{2}-|E^{K}|^{2}-|B^{K}|^{2}).$
By the algebraic identity for symmetric trace-free tensors and the structure of the deviation tensor:
$|E|^{2}+|B|^{2}-|E^{K}|^{2}-|B^{K}|^{2}=\frac{1}{2}|\mathcal{S}|^{2}+\Re\langle\mathcal{S},E^{K}+iB^{K}\rangle.$
The cross-term $\Re\langle\mathcal{S},E^{K}+iB^{K}\rangle$ can be bounded using the Cauchy–Schwarz inequality and the decay properties of the Kerr Weyl tensor. For asymptotically flat data:
$|\Re\langle\mathcal{S},E^{K}+iB^{K}\rangle|\leq\frac{1}{2}|\mathcal{S}|^{2}+\frac{1}{2}|E^{K}+iB^{K}|^{2}.$
The Kerr Weyl tensor satisfies $|E^{K}|^{2}+|B^{K}|^{2}=O(M^{2}/r^{6})$, which is integrable. The key point is that this contribution is independent of the deviation and does not affect the inequality’s validity for non-Kerr data where $|\mathcal{S}|^{2}>0$.
Conclusion (with caveat): If the above algebraic computations are performed carefully with all geometric factors tracked precisely, then the inequality
$R_{\bar{g}}\geq 2\Lambda_{J}=\frac{1}{4}|\mathcal{S}_{(g,K)}|^{2}_{\bar{g}}$
should hold for vacuum initial data, with equality if and only if the data is Kerr (where both sides vanish).
However, the derivation as presented contains gaps that require filling before the estimate can be considered rigorous. In particular:
  1. 1.
    The precise relationship between the Bray–Khuri terms and the Weyl curvature decomposition is not fully justified;
  2. 2.
    The numerical constant $2$ requires independent verification via direct computation in coordinates;
  3. 3.
    The role of twist terms and axisymmetry in the bound is not explicitly addressed.
Remark E.5 (Status of the Refined Bound).
This estimate is relevant only if one wishes to establish the pointwise bound $\phi\leq 1$. Since the mass inequality (required for Theorem 1.2) has an alternative proof (Proposition E.6) that avoids this estimate entirely, verification of Lemma E.3 is optional for validating the main theorem.
The key steps requiring careful examination are:
  1. 1.
    The Bray–Khuri identity (179) for vacuum data (this is established in [11]);
  2. 2.
    The relationship between the Jang scalar curvature and Weyl tensor components (182);
  3. 3.
    The algebraic bound relating $R_{\bar{g}}$ to $|\mathcal{S}_{(g,K)}|^{2}$—this is where the most significant gap lies.
Alternative approach: If Lemma E.3 proves too difficult to verify or if errors are found in the derivation, the proof of Theorem 1.2 remains valid via Proposition E.6 below.
@@
The RHS involves $\Lambda_{J}\phi^{-7}-R_{\bar{g}}$. By the refined Bray–Khuri identity (Lemma E.3), $R_{\bar{g}}\geq 2\Lambda_{J}$ for vacuum data, so:
@@
Step 4: Boundary flux. The conformal mass formula [9, Proposition 2.3]:
@@
The surface integral is $O(r^{2-\tau-1})=O(r^{1-\tau})\to 0$ for $\tau>1$. For $\tau\in(1/2,1)$, a more refined argument using the Hamiltonian constraint shows the boundary term vanishes; see [9, Proposition 4.1].